OpenStudy (anonymous):

What is the solution to the system below? -3x-y=-3 9x+3y=12 (3, 3) (3, -3) no solution infinite solutions on the line

OpenStudy (bahrom7893):

This is how you would solve it: -3x-y = -3 9x+3y=12 If I simplify the 2nd equation a bit, I will get: -3x - y = -3 3x + y = 4

OpenStudy (bahrom7893):

Now: Solve for 3x in the 2nd equation and plug it into the first one: -3x - y = -3 3x = 4 - y

OpenStudy (bahrom7893):

if you plug that into the first one, you will get: - ( 4-y ) - y = -3 -4 + y - y = -3 -4 = -3 No solutions exist.

OpenStudy (anonymous):

davette this is too long a solutions for this question better check the ratios a1 b1 c1 ---- ---- ---- a2 b2 c2 IF a1 b1 ---- not equal to ---- => one solution a2 b2 If a1/a2 = b1/b2 = c1/c2 => infinite solutions if a1/a2 = b1/b2 not equal to c1/c2 => no solution